OpenStudy (anonymous):

simplify x3-27 over x3+3x2+9x

OpenStudy (anonymous):

x3 means x cubed lucy?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i dont think this simplifies

OpenStudy (anonymous):

x^3-27 = (x-3)(x^2+3x+9) yes it does - sorry

OpenStudy (anonymous):

lol.... its fine

OpenStudy (anonymous):

(x-3)(x^2+3x+9) / x(x^2+3x+9) = (x-3)/x

OpenStudy (anonymous):

ok?